### v1

# Repurposing 2D Diffusion Models with Gaussian Atlas for 3D Generation

Abstract
Recent advances in text-to-image diffusion models have been driven by the increasing availability of paired 2D data. However, the development of 3D diffusion models has been hindered by the scarcity of high-quality 3D data, resulting in less competitive performance compared to their 2D counterparts. To address this challenge, we propose repurposing pre-trained 2D diffusion models for 3D object generation. We introduce Gaussian Atlas, a novel representation that utilizes dense 2D grids, enabling the fine-tuning of 2D diffusion models to generate 3D Gaussians. Our approach demonstrates successful transfer learning from a pre-trained 2D diffusion model to a 2D manifold flattened from 3D structures. To support model training, we compile GaussianVerse, a large-scale dataset comprising 205K high-quality 3D Gaussian fittings of various 3D objects. Our experimental results show that text-to-image diffusion models can be effectively adapted for 3D content generation, bridging the gap between 2D and 3D modeling.
1 Introduction
Understanding the 3D world is crucial for numerous real-world applications. In this work, we focus on generative 3D modeling with the primary objective of generating high-quality 3D assets from given textural descriptions [13, 32].
Figure 1: Previous 3D generators are either end-to-end 3D models [55, 10] or a combination of a multi-view 2D generator followed by a 2D-to-3D lifting model [52, 48]. In this work, we achieve 3D object generation by directly fine-tuning 2D generation models.
With advances in 3D content representation, diffusion models [11, 43] begin to play a predominant role in 3D generation [24, 30, 12]. Among different 3D representations, 3D Gaussians [15] have gained popularity for modeling 3D scenes and objects due to their high efficiency, explicit structure, and superior rendering quality. Notably, training diffusion models to generate 3D Gaussians requires high-quality “ground truth” 3D Gaussians, which must be pre-fitted to the 3D objects [10, 45, 26].
To advance the community, we constructed a large-scale dataset named GaussianVerse, comprising high-quality 3D Gaussian fittings for a diverse range of objects. Unlike previous studies that also pre-compute 3D Gaussian references, GaussianVerse provides higher-quality fittings with the minimum required number of Gaussians, through a novel pruning strategy and a more effective, but compute-intensive, fitting process.
Designing standalone diffusion models for 3D generation is straightforward. However, such models have significant limitations especially when trained solely on 3D data, as high-quality 3D data is relatively scarce compared to 2D images. To achieve higher rendering fidelity in 3D generation, it is desirable to leverage the learned prior knowledge from well-pretrained 2D diffusion models [32, 5, 20, 27]. Some methods have been proposed to incorporate pretrained 2D diffusion models with frozen weights into the 3D generation pipeline through complex designs, such as score-distilled sampling [32] and collaborative controls [5].
In this paper, we propose a fresh perspective that repurposes 2D diffusion models for 3D generation through direct fine-tuning. To fully harness the capabilities of these 2D diffusion models, we introduce Gaussian Atlas, a novel 2D representation of 3D Gaussians. This representation enables direct fine-tuning by first transporting unorganized 3D Gaussians into a standard 3D sphere and then applying equirectangular projection to map them into a square 2D grid, creating a Gaussian Atlas. We show that these Gaussian atlases facilitate transfer of the prior knowledge captured in 2D diffusion models to the 3D generation task. By doing so, our approach provides a means to leverage the learned 2D priors for 3D generation, unlocking new possibilities for efficient and effective 3D content creation.
To summarize, our major contributions are three-fold: (i) We present a large-scale dataset, GaussianVerse, consisting of 205,737 high-quality 3D Gaussian fittings for diverse objects sampled from Sketchfab [42]; (ii) We propose a novel 2D representation of 3D Gaussians, Gaussian Atlas, facilitating 3D generation tasks through a new perspective by fine-tuning pretrained 2D diffusion models; (iii) We demonstrate that the proposed approaches surpass state-of-the-art 3D Gaussian generators in terms of both generation-prompt alignment and user preferences.
2 Related Work
Repurposing 2D diffusion models. Diffusion models [11, 36, 6], are originally designed for text-to-image generation. When trained on large-scale internet images, prior knowledge learned from diverse domains can be transferred to different tasks. Marigold [14] is an example of how fine-tuning pretrained Latent Diffusion (LD) models [36] can benefit depth prediction. Marigold first re-parameterizes single-channel depth maps as RGB images and uses the pretrained VAE encoder to compress both the input image as well as its paired depth map into a shared latent space. The two latents are then concatenated to enable conditional generation of depth latents, which are subsequently decoded back into RGB space through the VAE decoder. More recent work, such as GeoWizard [7] and GenPercept [50], explored the repurposing of LD models for more geometric prediction tasks and beyond. As more and more studies have shown the potential of finetuning LD for different purposes, this work is among the first to explore how a LD model can be adapted for 3D generation, addressing the significant disparity between 3D objects and 2D images.
Figure 2: (a) GaussianVerse offers high-quality 3DGS fittings for diverse 3D objects. (b) A word cloud and (c) a list of the most frequently occurring words generated from the captions [25] of the fitted objects. (d) The distribution of the number of Gaussians per fitting shows that our method generates fewer Gaussians for simple objects and more for complex ones. Red dotted lines indicate the 25%, 50%, and 75% percentile of the distribution.
3D generation with Gaussian splatting. Unlike 2D images, 3D objects are more difficult to generate due to the additional dimension and geometric constraints [28, 15]. Among all 3D generation methods [47, 34, 16, 4, 48, 51, 18], there is a line of work that utilizes diffusion models to generate 3D Gaussians, which is closely related to this paper. GSD [29] introduced rendering guidance to constrain the sampling of 3D Gaussians with 2D observations. L3DG [35] is another example, which embeds 3D Gaussians onto a dense latent grid and learns a diffusion model in the latent space for generation. Moreover, GVGen [10] and GaussianCube [55] achieve 3D Gaussian diffusion by transforming sparsely located Gaussians into more structured 3D volumes. GVGen fits 3D Gaussians directly on a volume with offsets, while GaussianCube applies Optimal Transport to move unorganized 3D Gaussians to vertices of a predefined 3D grid. Most recently, Gaussian Anything [17] adopts a two-step approach that generates Gaussians by auto-encoding latents in the point cloud space. In this work, we propose a novel structuralization approach that enables 2D transformation for 3D Gaussians, to leverage the power of well pretrained 2D diffusion models and unify the architecture of both 2D and 3D generation.
2D representations of 3D content. Representing 3D objects as 2D planes is not new. A major line of research encodes 3D structures using implicit Triplanes [3]. NFD [41] was one of the first attempts to train 2D diffusion models to generate Triplanes, from which neural fields can be derived to reconstruct 3D objects. To enhance 3D coherence, CRM [48] integrates multi-view observations into the Triplane generation process. InstantMesh [51] and Instant3D [18] follow a similar design but employ more sophisticated models to improve Triplane synthesis. Instead of generating neural fields, TriplaneGaussian [58] decodes 3D Gaussian attributes directly from generated Triplanes. More recently, DiffGS [57] has refined the TriplaneGaussian paradigm by learning a mapping between pre-fitted 3D Gaussians and Triplanes via a latent diffusion model. Several works share motivations similar to ours. For example, PI3D [20] fine-tunes a pretrained text-to-image generation model to produce Triplanes represented as “pseudo-images,” while HexaGen3D [27] leverages pretrained image diffusion models for improved Triplane generation. However, their fine-tuned diffusion models exhibit limited capacity, and an additional minutes-long refinement process is typically required to enhance the 3D generation.
In addition to Triplane-based representations, Splatter Image [44] trains a 2D network to infer pixel-wise Gaussian attributes from single images; Omages [53] trains 2D diffusion models to fit 2D UV maps of the geometry and materials of 3D objects, ultimately reconstructing a textured 3D mesh. However, Omegas supports only class-conditioned generations of a limited pre-defined categories; GIMDiffusion [5] adopts a similar approach by representing the surface of 3D objects with geometry images [8] and leveraging 2D diffusion models in the UV space. However, their generated UV maps are sparse, leading to artifacts at the seams of the assembled 3D mesh. In contrast, our method introduces a dense 2D representation of 3D Gaussians, thereby mitigating these limitations.
3 GaussianVerse
In this section, we present GaussianVerse, a large-scale dataset containing high-quality 3D Gaussian fittings for a wide range of 3D objects. A summary of this dataset is available in Table 1 and Figure 2. We describe the process of fitting per-object 3D Gaussians below. More details are provided in the supplementary materials.
# Gaussians per fitting Total # fittings Compute time
10,435$\pm$4,453 205,737 $>$3.8 GPU Years
Table 1: Statistics for the proposed GaussianVerse dataset. We report the average (with standard deviation) number of 3D Gaussians per fitting, the total number of fitted objects, and the overall compute time spent on A100 GPUs.
Preliminaries: 3D Gaussian splatting [15]. Given multi-view observations of a 3D object, 3D Gaussian Splatting (3DGS) represents the scene as a set of Gaussian distributions defined in 3D space. Each Gaussian is characterized by five attributes: its 3D location $\mathbf{x_{i}}\in\mathcal{R}^{3}$, color information $\mathbf{c_{i}}\in\mathcal{R}^{3}$, opacity $\mathbf{o}_{i}\in\mathcal{R}^{1}$, scale $\mathbf{s}_{i}\in\mathcal{R}^{3}$, and 3D rotation $\mathbf{r}_{i}\in\mathcal{R}^{3}$. The 3D covariance matrix $\Sigma_{i}$ can be represented as $\Sigma_{i}=\mathbf{r}_{i}\mathbf{s}_{i}\mathbf{r}_{i}^{T}$. A 2D image $C$ can be rendered from properly structured 3D Gaussians through $\alpha$-blending:
$C^{\pi}=\sum_{j=1}\mathbf{c}_{j}\omega_{j}^{\pi}\prod_{k=1}^{j-1}(1-\omega_{k}^{\pi}),$ (1)
where $\pi$ is the camera pose, $\omega_{i}^{\pi}$ represents $z$-depth ordered opacity, computed as $\mathbf{o}_{i}e^{-\frac{1}{2}(\mathbf{x}-\mathbf{x}^{\pi}_{i})^{T}(\Sigma_{i}^{\pi})^{-1}(\mathbf{x}-\mathbf{x}^{\pi}_{i})}$. The terms $(\mathbf{x}_{i}^{\pi},\Sigma_{i}^{\pi})$ denote the 2D Gaussian projection on the image plane derived from the 3D Gaussians with $\pi$.
Figure 3: Representing 3D Gaussians as 2D Gaussian Atlas. For each fitting, 3D Gaussians are first translated to the surface of a standard sphere $S$ via optimal transport. Then, the surface Gaussians are flattened onto the 2D plane via equirectangular projection with reusable indices. We obtain Gaussian Atlas by reorganizing the flattened coordinates to pixels of a dense 2D square of size $\sqrt{N}\times\sqrt{N}$.
3D Gaussian fitting with bounded quantity. For training diffusion models, high-quality 3D Gaussian “ground truth” need to be pre-fitted for 3D objects in different formats. We build our fitting model upon the state-of-the-art, Scaffold-GS [23], along with non-trivial modifications. First, we exclude view properties from the MLP predictors for attribute querying to enable more view-invariant applications. Second, we propose constraining the number of Gaussians per fitting, as also suggested in [29, 55].
Different 3D objects can exhibit diverse geometries and appearances, leading to highly variable 3DGS fittings. A common approach is to have a constant number of Gaussians per fitting [29, 55], which enforces a uniform distribution of number of Gaussians across all 3D objects. Consequently, low-poly, simple objects can get excessively parameterized, with low, though not zero, scale and opacity values. Instead of constraining the number of Gaussians to be constant, we introduce the visibility ranking strategy that bounds the number of fitted Gaussians at $\tau$, a self-defined parameter. Specifically, during 3DGS fitting, we track each Gaussian’s opacity to assess its visibility from randomly selected camera views. If the number of Gaussians exceeds the bound $\tau$ after densification, we sort their opacities and prune those with the lowest values. This pruning strategy differs from the original implementation [15], which only prunes completely invisible Gaussians. Additionally, to align with Scaffold-GS’s design, we apply the same ranking and pruning based on the tracked visibility of anchors. Final Gaussians for pruning are selected by averaging the ranks from both tracking metrics.
Our visibility ranking strategy introduces minimal computational overhead while providing significant flexibility for various applications. See Figure 11 in the supplementary materials for our fitting results on objects with varying levels of detail and Figure 2 (d) for the distribution of the number of Gaussians across all fittings in GaussianVerse.
With the above fitting strategy, we optimize per-object 3D Gaussians by minimizing photometric losses against multi-view RGB renderings:
$\lambda_{rgb}^{\prime}\mathcal{L}_{rgb}+\lambda_{ssim}^{\prime}\mathcal{L}_{ssim}+\lambda_{lpips}^{\prime}\mathcal{L}_{lpips}+\lambda_{reg}^{\prime}\mathcal{R},$ (2)
where $\mathcal{L}_{rgb}$ represents the color space L1 loss, $\mathcal{L}_{ssim}$ is the negated structural similarity index, $\mathcal{L}_{lpips}$ is the perceptual loss, and $\mathcal{R}$ is a scaling regularization term [23]. The $\lambda$’s are weights for each loss. Note that the addition of perceptual loss nearly doubles the training time, but it is necessary for high rendering fidelity with fewer Gaussians.
Dataset details. GaussianVerse consists of 205,737 3DGS fittings of 3D objects sampled from Sketchfab [42] under CC licenses. We used the multi-view 2D renderings provided in [59] to perform per-object fittings. The 2D renderings came in 10 different categories, and we used renderings from 9 of the categories, excluding those of ’poor quality’. This curated list results in a wide coverage of diverse objects. We visualize a word cloud as well as the top 20 most frequent words from the captions [25] in Figure 2 (b) and (c). For optimal efficiency and quality balance [55, 10], we set the bound $\tau$ to $192\times 192=36,864$. A cluster of A100 GPUs is employed for large-scale parallel fitting. Each 3DGS fitting job converges at around 20,000 steps, translating to approximately 10 minutes of fitting time per object, with a total of over 3.8 A100 GPU years spent.
Comparison with related work. Compared with several previous studies [29, 10, 55] which also fit per-object 3D Gaussians for training diffusion models, we achieve higher-quality 3DGS fittings with fewer valid Gaussians but increased compute costs with a refined pruning strategy and a more effective, but compute-intensive, fitting process.
Figure 4: Repurposing 2D diffusion models for 3D Gaussian generation. Our pipeline consists of two stages. In the 3DGS pre-fitting stage (section 3), we pre-fit high quality 3D Gaussians for a diverse array of 3D objects with multi-view observations. The large-scale 3DGS fittings are then re-structured as Gaussian Atlas (section 4) with 2D representations for each 3DGS attribute. In the diffusion model training stage (section 5), we leverage the transformed 2D Gaussian atlases to repurpose a pretrained latent diffusion model (the 2D UNet denoiser $\mathcal{F}$) and ultimately achieve 3D content generation.
4 Formulating 3D Gaussians as 2D Atlas
In this section, we introduce a novel approach that transforms unorganized Gaussians in the 3D space to a dense 2D representation, namely Gaussian Atlas, making it possible to repurpose 2D diffusion models, for example, Latent Diffusion (LD) [36], for 3D generation tasks.
Motivation. LD can understand complex natural language and generate coherent 2D images, benefiting from the vast availability of over billions of paired text-image data [39]. However, text-to-3D generation presents greater challenges due to two key reasons: (i) the scarcity of large-scale datasets with 3D models comparable to those in 2D, as creating and annotating high-quality, textured 3D models remains both resource-intensive and time-consuming [25]; and (ii) the inherently higher-dimensional nature of 3D object representations, which impose complex geometric constraints, making it more difficult for diffusion models to interpret. These disparities in data availability and representational complexity between 2D and 3D motivate our approach to leveraging learned priors from pre-trained 2D diffusion models for 3D Gaussian generation.
However, sparse and unstructured Gaussians in 3D space cannot be directly passed to 2D models, which require inputs $\mathbf{X}$ to have: (i) only 2 spatial dimensions $W$ and $H$ such that $\mathbf{X}\in\mathbb{R}^{W\times H\times\cdot}$; (ii) valid “pixels” at each vertex of a dense 2D grid; and (iii) values within a specific distribution, either $\mathbf{X}\in[-1,1]$ for the VAE or $\mathbf{X}\sim\mathcal{N}(0,1)$ for the latent UNet. To make 3D Gaussians compatible with 2D diffusion models, we propose Gaussian Atlas, a 2D representation of 3D Gaussians.
One simple idea of 2D transformation is to project 3DGS onto an image plane with their 3D coordinates and given camera parameters. However, this naive approach loses depth information entirely and disrupts the original 3D structure’s topology, which is essential for 3D generation as it relies on accurate representation of 3D continuity. This highlights the need for a method that not only maps 3D Gaussians onto a 2D plane but also preserves 3D continuity to some extent. Given these requirements, we argue that methods similar to UV texture unwrapping, which unfold the surface of a 3D geometry onto a 2D plane, are more suitable. However, UV maps are usually not universally applicable since careful designs are needed for different 3D geometries [1, 9]. Therefore, instead of exact mappings between 3D geometries and their 2D UVs, we focus on one characteristic of UV unwrapping — geometry flattening.
Specifically, our goal is to find a mapping function $\mathcal{M}(\cdot)$ that transforms the 3D positions $\{\mathbf{x}\in\mathbb{R}^{3}\}$ of 3D Gaussians to 2D planar coordinates $\{\mathbf{\hat{x}}\in\mathbb{R}^{2}\}$: ($\mathcal{M}(\{\mathbf{x}\})\rightarrow\{\mathbf{\hat{x}}\}$. A natural approach is to parameterize such a function $\mathcal{M}$ as neural networks which can be optimized toward 2D and 3D consistency [56]. However, this approach requires repeated training of $\mathcal{M}$ on different objects which is not only time consuming but also makes the mapping process inconsistent between objects given that the heuristics for 2D flattening may be different for different geometries. As a result, diffusion models are not able to capture the irregular patterns and fail to generate meaningful contents. Experiments are provided in section 6.2 to support this claim. To this end, we seek a more coherent transformation between 3D and 2D with a simple and deterministic mapping function.
We formulate our transformation process into three sequential stages, as outlined in Figure 3. In the first stage, we hypothesize a unit sphere $\mathcal{S}$, which is centered at the origin of 3D-axis with unit radius, to be parameterized by $N$ 3D points $\{s_{i}\in\mathbb{R}^{3}\}$ that are uniformly distributed on its surface. Sphere is a well-studied structure that offers multiple standards for 2D projection, which makes it suitable for being considered in the flattening process. We then translate the unorganized 3D Gaussians to the surface points $\{s_{i}\in\mathbb{R}^{3}\}$ of $\mathcal{S}$ using Optimal Transport (OT) [2]. We call this process sphere offsetting. Our method differs from [55] by offsetting 3D Gaussians directly onto the surface, rather than vertices within a volume. Our transportation process is also much faster due to the adaptive number of Gaussians.
After positioning the 3D Gaussians on the surface of a unit sphere, we employ the classic equirectangular projection [49] as our function $\mathcal{M}$ to obtain flattened 2D coordinates $\{p_{i}\in\mathbb{R}^{2}\}$ of the 3D Gaussians. Lastly, in the plane offsetting stage, we apply another OT to map $\{p_{i}\}$ to the vertices $\{q_{i}\in\mathbb{R}^{2}\}$ of a square 2D grid with spatial size $\sqrt{N}\times\sqrt{N}$, further reducing sparsity. In particular, since now $\mathcal{M}$ is a deterministic function, the mapping between $\{p_{i}\}$ and $\{q_{i}\}$ remains identical between different 3D objects. This consistency allows us to perform OT in the last stage only once and reuse the computed indices for all objects. We refer to the final grid-like 2D representation of 3D Gaussians as Gaussian Atlas, with each atlas $\mathbf{X}$ in the shape of $\sqrt{N}\times\sqrt{N}\times(||\mathbf{x}-\mathbf{s}||+||\mathbf{c}||+||\mathbf{o}||+||\mathbf{s}||+||\mathbf{r}||)$ entailing all attributes of 3D Gaussians.
5 2D Diffusion for 3D Gaussian Generation
After transforming fitted 3D Gaussians onto 2D planes, we are able to fine-tune the pre-trained Latent Diffusion (LD) [36] with Gaussian atlases.
Preliminaries: Latent Diffusion. LD is a family of models that generate 2D images based on text prompts. The core components of LD include a Variational AutoEncoder (VAE) and a UNet $\mathcal{F}(\cdot)$. The VAE encoder compresses images into lower-dimensional latents $l$ to facilitate efficient diffusion in the latent space. By injecting Gaussian noise to the latents, $\mathcal{F}$ can be trained through self-supervised denoising via v-parameterization [38]:
$\mathcal{L}_{diff}=\mathbb{E}_{l_{0},\mathbf{z},t}\left[\left\|\nabla_{l_{t}}\mathbf{z}-\nabla_{l_{t}}\mathcal{F}(l_{t},t)\right\|^{2}\right],$ (3)
where $l_{t}$ is the noisy latent at timestamp $t$, $\nabla_{l_{t}}\mathbf{z}$ is the ‘velocity’ added to $l_{t}$. During inference, a sample can be generated through the reverse diffusion process by iteratively forwarding $\mathcal{F}$ for denoising. The VAE decoder then upsamples the generated latent back to the original RGB space.
Finetuning LD with Gaussian Atlases. The standard fine-tuning approach for LDs involves VAE-based encoding and decoding [14]. However, in the supplementary materials, we argue that such auto-encoding is unsuitable for Gaussian atlases - the Gaussian attributes differ from the natural images used to train the LD. Finetuning LD UNet with 2D atlases is feasible only if the distributions of the atlases are aligned with the VAE-encoded latents. We therefore standardize the 2D atlases using the pixel-wise mean and standard deviation computed from the entire GaussianVerse. This normalization prevents extreme high- or low-frequency pixel values while preserving the characteristics of the original VAE-encoded image latents, ultimately accelerating the fine-tuning process.
Figure 5: Qualitative Comparisons. Our 3D generations exhibit the highest quality, minimal artifacts, and the best alignment with text prompts. In contrast, DreamGaussian [45], LGM [46], and TriplaneGaussian [58] struggle to produce natural and complete 3D assets, while GaussianCube [55] fails to capture key phrases in text prompts, such as “on a plate” and “pink rooftop.”
3D Gaussians are characterized by five attributes: the 3D location (mean) $\mathbf{x}$, albedo $\mathbf{c}$, and scale $\mathbf{s}$, each represented as three-channel features; opacity $\mathbf{o}$, which is a single-channel feature; and rotation $\mathbf{r}$, typically represented as a four-channel quaternion. Since the LD UNet was originally trained on four-channel latents, we pad each three-channel attribute with the one-channel opacity and repeat the input layer of the UNet four times [14, 7] to accommodate all attributes as inputs. The final Gaussian atlases, used for fine-tuning the UNet, have a shape of $128\times 128\times 16$. During rendering, we appropriately unpack the different attributes and average the three repetitions of opacity.
We finetune the LD UNet $\mathcal{F}$ with a combination of diffusion loss (Equation 3) and photometric loss between the renderings from denoised Gaussians and the reference images. The final objective for tuning $\mathcal{F}$ becomes:
$\lambda_{diff}\mathcal{L}_{diff}+\lambda_{rgb}\mathcal{L}_{rgb}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{lpips}\mathcal{L}_{lpips},$ (4)
where $\mathcal{L}_{rgb},\mathcal{L}_{mask}$ are L1 loss on RGB renderings and accumulated opacity maps and $\mathcal{L}_{lpips}$ is the perceptual loss. $\lambda$s are the corresponding loss weights. The overall pipeline is outlined in Figure 4.
6 Experiments
In this section, we benchmark and evaluate the effectiveness of our proposed methods on zero-shot text-to-3D generation, one of the most fundamental tasks in generative 3D modeling.
Figure 6: Additional qualitative results. Our method effectively repurposes 2D diffusion models for high-quality 3D contents. The generated Gaussian atlases are presented in the order from top left to bottom right: 3D location $\mathbf{x}$, albedo $\mathbf{c}$, color-coded opacity $\mathbf{o}$, normalized scale $\mathbf{s}$, and the last three channels of normalized quaternion $\mathbf{r}$. More results can be found in the supplementary materials.
Implementation Details. Given the upper bound $\tau=36,864$ of valid 3D Gaussians during the pre-fitting stage, we set the total number of 3D Gaussians per Gaussian atlas to $N=16,384=128\times 128$. This choice of $N$ accommodates the large variance of 3D Gaussians in GaussianVerse.
We initialize the 2D UNet $\mathcal{F}$ from a LD checkpoint pretrained on LAION [39] and use CLIP [33] to encode text prompts. The UNet is then fine-tuned with classifier-free guidance by randomly zeroing out the text encodings with a probability of 20%. We employ the AdamW optimizer [21] with a learning rate of $5\times 10^{-5}$ and a batch size of 64. Exponential Moving Average (EMA) and mixed precision training are enabled for stable and efficient training. For fair comparisons with the baseline method [55], only approximately 130K atlases from GaussianVerse were used for training. The paired text prompts were from [25] and the multi-view references were from [59]. The LD UNet was finetuned for 1M steps on 8 A100 GPUs.
For inference, we start from 2D random noise in the same shape as a Gaussian atlas and generate clean samples via reverse diffusion using the adaptive DPMsolver+++++ + [22] with a guidance scale of 3.5. A 3DGS sample can be generated and rendered in less than 5 seconds.
Metrics. Reference based metrics are not feasible for evaluating zero-shot text-to-3D generations. Following [32, 13, 55], we evaluate the alignment of text prompts and 10 random 2D renderings of the 3D generations with CLIP score [33] as well as VQA score [19] on 100 diverse test prompts. Additional user studies and qualitative comparisons were also conducted for comprehensive evaluations.
Comparisons. We compare our proposed method against four representative methods in the field of 3D Gaussian generation: An optimization based method — DreamGaussian [45]; A method uses also a 2D model — LGM [46]; A method generates also 2D representations — TriplaneGaussian [58]; and the state-of-the-art 3D Gaussian generator with a 3D diffusion model — GaussianCube [55]. For LGM and TriplaneGaussian, we utilize MVDream [40] to generate 2D images from text prompts to initiate 3D generations. Notably, since there are no open-source implementations of the relevant works that also adopt 2D representations of 3D content [20, 27, 5], direct comparisons are not possible.
Methods CLIP score $\uparrow$ VQA score $\uparrow$ # Gaussians $\downarrow$
DreamGaussian [45] 20.52 0.37 40K
LGM [46] 20.28 0.35 66K
TriplaneGaussian [58] 21.10 0.46 16K
GaussianCube [55] 22.31 0.52 33K
GaussianAtlas (Ours) 23.20 0.61 16K
Table 2: Qualitative comparisons. Our method achieves performance comparable to the state-of-the-art in terms of CLIP similarity scores, with the minimum number of 3D Gaussians.
6.1 Results
Qualitative results. We present zero-shot 3D generation results in Figure 5. Compared to DreamGaussian and LGM, our generations exhibit high visual fidelity without excessive details. In contrast to TriplaneGaussian and GaussianCube, our method aligns closely with the text prompts, reflecting more precise prompt conditioning in the generations. According to the additional qualitative results shown in Figure 6, our fine-tuned 2D diffusion model is capable of generating high-quality Gaussian atlases that effectively generalize to 3D information. Consequently, our generated 3D contents are coherent and self-contained, with no seams or artifacts that would typically be observed from UV unwrapping methods. We provide more qualitative results in the supplementary materials.
User studies [^1] The user study was conducted entirely by researchers at Stanford.. The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts. They performed pairwise comparisons between generations by our method and those from a competing method. We calculated the win rate from the collected $>2,500$ valid responses. As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube. The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
User studies. The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts. They performed pairwise comparisons between generations by our method and those from a competing method. We calculated the win rate from the collected $>2,500$ valid responses. As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube. The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
Figure 7: User study results. Our method outperforms state-of-the-art methods [55, 58] in user preferences regarding generation quality and alignment with text prompts.
6.2 Discussions and Ablations
This work takes a step toward unifying 2D and 3D generation by transforming 3D Gaussians onto 2D planes, allowing the direct fine-tuning of a pre-trained 2D diffusion model for generating 3D contents. In this section, we conduct extensive studies and ablative experiments to validate our proposed approach from three key perspectives.
$\bullet$  Learned prior knowledge in text-to-image diffusion models is universally transferable. Without pre-training on large-scale 2D datasets, the model struggles to develop a comprehensive understanding of natural language and content generation when trained solely on 3D data. As shown in Figure 8, we effectively repurpose the pre-trained 2D diffusion model, resulting in significantly faster convergence and higher-quality 3D generation. The qualitative results show that, compared to training from scratch, our approach produces more coherent and structured 3D assets earlier in training. The quantitative results further support this finding, as our repurposed models consistently achieve higher scores in different training steps.
Methods # Training steps CLIP score $\uparrow$ VQA score $\uparrow$
Train from scratch 500K 19.33 0.23
Repurpose 2D LD 500K 21.61 0.49
Train from scratch 1M 20.85 0.40
Repurpose 2D LD 1M 23.20 0.61
Figure 8: Finetuning from a pretrained 2D diffusion model leads to faster generalization. Top: 3D generations at different training checkpoints from finetuning (top row) and training from scratch (bottom row) using the same prompt ‘A toy hamburger with beef patty & cheese’. Bottom: Quantitative comparisons at checkpoints with different training steps.
$\bullet$  Gaussian atlases yield consistent and capturable visual patterns, which makes the pre-trained 2D diffusion model easier to generalize towards. To validate the claim, we ablate on the “flattening” strategy and adopted an optimization based approach [56] to find the most natural cut for each object and perform 3D-to-2D transformation via a combination of neural networks. We train the networks on a per-object basis for the entire GaussianVerse, which roughly took another A100 GPU year. The flattened 3D Gaussians are then used to finetune the LD UNet by following the same training procedure. We visualize generation results in Figure 9. With capturable visual patterns, our proposed flattening strategy leads to faster model convergence and much higher generation quality at the same training step.
Figure 9: The optimization-based flattening approach results in inconsistent patterns and poses challenges for fine-tuning. We present two objects (a) and their corresponding Gaussian atlases (3D location $\mathbf{x}$ only) obtained using our proposed flattening approach (b) and the optimization-based approach [56] (c), which generates only noisy Gaussians (d) and atlases (e) after training for the same number of steps.
$\bullet$  A pre-trained 2D diffusion model serves as a better initialization for 3D generation. In Figure 10, we plot the weight differences between our fine-tuned UNet and both a randomly initialized UNet and the pre-trained UNet directly from LD. We observe overall small changes in the weights from the pre-trained UNet. Even at the layer with the greatest change, the difference is still 8×8$\times$8 × smaller than that from a random initialization. This indicates that the pre-trained LD weights provide better starting points for 3D generation, facilitating easier convergence to a local minimum.
Figure 10: Log scale mean squared errors (smoothed) of per layer weights between different UNets. Finetuning from a pre-trained LD UNet leads to smaller weight differences.
7 Conclusion
This paper presents an advancement in unifying 2D and 3D content generation by repurposing a pre-trained 2D diffusion models to generate 3D Gaussians. Historically, fine-tuning 2D diffusion models with 3D Gaussians was infeasible. We introduce Gaussian Atlas, a novel approach that transforms unorganized 3D Gaussians into a structured 2D grid, thereby enabling the effective utilization of 2D networks for 3D generation. We also present GaussianVerse, a large-scale dataset comprising 205,737 3D Gaussian fittings of diverse 3D objects, which facilitates diffusion model training. Experiments demonstrate that our method achieves state-of-the-art 3D generation results using significantly fewer Gaussians. We also provide extensive discussions and ablation studies that validate our claims and offer valuable insights into the effectiveness of our approach.
Supplementary Material
Appendix A Supplementary for GaussianVerse
Methods Venue # of 3DGS fittings # of Gaussians per fitting
GSD [29] ECCV 2024 6,000 1,024
GVGen [10] ECCV 2024 $\sim$46,000 32,768
GaussianCube [55] NeurIPS 2024 125,653 32,768
ShapeSplat [26] 3DV 2025 $\sim$65,000 $>$20,000
GaussianVerse (ours) - 205,737 10,435
Table 3: GaussianVerse contains large-scale 3DGS fittings with adaptive number of Gaussians, which allows various applications.
A.1 Fitting Details
As described in Section 3, we adopted Scaffold-GS [23] as the base model for per-object 3D Gaussian fitting. We observed duplicated Gaussians when the number of offsets was large, even at the default value. Therefore, we reduced the number to 4 to allow more anchors to be initialized with random values. The 3D Gaussians were then optimized according to the objective in Equation 2, with $\lambda_{rgb}^{\prime}$ set to 0.8, $\lambda_{ssim}^{\prime}$ set to 0.2, $\lambda_{lpips}^{\prime}$ set to 0.02, and $\lambda_{reg}^{\prime}$ set to 0.01.
A.2 3DGS Quality of GaussianVerse
We compare our proposed 3DGS fitting method in Section 3 with the densification-constrained fitting method introduced by GaussianCube [55] and the upper-bound method, Scaffold-GS [23], in Figure 11. Compared to GaussianCube, our method achieves better rendering quality. Compared to the upper bound, our results show no significant visual degradation while using significantly fewer 3D Gaussians. As shown in Figure 11, the rendering quality of fitted 3DGS does not improve with an increased number of Gaussians. This observation confirms that many Gaussians contribute minimally to the final rendering quality and can be merged or pruned during fitting.
A.3 Comparison with Other 3D Gaussian Datasets
GaussianVerse is a large-scale dataset consisting of high-quality 3D Gaussian fittings for a wide range of objects. We note that there are a few other studies that also fit per-object 3D Gaussians to assist in the training of diffusion models. A direct comparison is provided in Table 3.
Appendix B Supplementary for Gaussian Atlas Formulation
B.1 Formulation Details
As discussed in Section 4, the adaptive nature of 3DGS fittings in GaussianVerse enables faster transformation of 3DGS to 2D Gaussian Atlas compared to the similar process in [55]. Notably, the computation time is significantly reduced during the non-square Optimal Transport step for sphere offsetting, since the number of 3D positions $\{\mathbf{x}\}$ is typically much smaller than $N$, the number of surface points $\{s\}$ on the standard sphere $S$. After sphere offsetting, for 3DGS fittings that have a smaller size than $N$, we pad additional Gaussians by duplicating the ones with the smallest scales and set their opacity to 0. For 3DGS fittings that have more Gaussians than $N$, we prune the Gaussians with the smallest scales before sphere offsetting. In practice, constructing one Gaussian atlas takes an average of approximately one minute.
Figure 11: Comparisons of 3DGS fitting methods. Our method achieves high fitting quality comparable to the upper bound with significantly fewer valid 3D Gaussians with positive opacity.
Appendix C Supplementary for Finetuning LD with Gaussian Atlas
C.1 Training Diffusion Model Without VAEs
The typical fine-tuning approach involves VAE encoding and decoding [14]. However, we argue that such a VAE auto-encoding is inappropriate for Gaussian atlases due to three reasons: (i) VAE encoding is a lossy compression of the atlases, whose accuracy is crucial for Gaussian rendering. We demonstrate the impact of using VAE for auto-encoding Gaussian atlases in Figure 12; (ii) Even at the maximum bound $\tau=36,864=192\times 192$, the number of Gaussians is significantly smaller than the number of pixels required for VAE input (e.g., $768\times 768$). Naive up-sampling of the Gaussian Atlas $\mathbf{X}$ would significantly increase computational costs; (iii) Even with scaling and shifting, $\mathbf{X}$ does not visually resemble the natural RGB images originally used for VAE training, but rather analogizes latent features. Considering these, we remove the original VAE and fine-tune the LD UNet directly on Gaussian atlases. We make appropriate modifications to the input and output layers of the UNet to accommodate all Gaussian attributes with the correct number of channels [14].
Figure 12: Latent diffusion VAE degrades 3D Gaussian quality. In the bottom row, we present a transformed Gaussian atlas and its reconstruction with and without VAE auto-encoding. When the atlas is up-sampled to a higher resolution, no obvious differences are observed in the atlas. However, after rendering the decoded Gaussians to images, we can observe much higher disparities, resulting in significantly reduced PSNR and visual quality. This suggests that incorporating the VAE for 3DGS diffusion is impractical.
C.2 More Details
Diffusion model training.
We set $\lambda_{diff}$ to 1.0, $\lambda_{rgb}$ to 10.0, $\lambda_{mask}$ to 1.0, and $\lambda_{lpips}$ to 1.0 to finetune the LD UNet. At each training step, we randomly sample one view of 2D renderings to compute the photometric losses. We enable gradient checkpoint to save GPU memory, which allows a local batch size of 8 for each GPU, with a total batch size of 64 across 8 GPUs.
Appendix D More Results
D.1 More Comparisons
Since Omages [53] does not support text-to-3D generation and the implementations of PI3D [20], HexaGen3D [27], and GIMDiffusion [5] are not publicly available, we provide additional qualitative comparisons against Splatter Image [44], a method that also generates 2D representations of 3D objects. As shown in Figure 13, while Splatter Image reconstructs the complete geometry of 3D objects with few artifacts, it struggles to generate coherent appearances with meaningful details. In contrast, our method leverages the prior knowledge embedded in a pre-trained 2D diffusion model, thereby yielding significantly improved generation quality.
Figure 13: Comparison with a 2D generation approach [44]. Note that we provide MVDream [40] generated 2D images (shown as the main images) to initialize 3D generations of [44] (shown as the smaller images).
Figure 14: Diverse text-to-3D generation results from the same prompt. We present the generated Gaussian atlases in the order from top left to bottom right: 3D location $\mathbf{x}$, albedo $\mathbf{c}$, color-coded opacity $\mathbf{o}$, normalized scale $\mathbf{s}$, and the last three channels of normalized quaternion $\mathbf{r}$.
Figure 15: Similar to the original LD, our repurposed LD is not good at capturing long text prompts for 3D generations.
Figure 16: More text-to-3D generation results. Our model is able to generate 2D Gaussian atlases to resemble high-quality 3D Gaussians from various prompts. We present the generated Gaussian atlases in the order from top left to bottom right: 3D location $\mathbf{x}$, albedo $\mathbf{c}$, color-coded opacity $\mathbf{o}$, normalized scale $\mathbf{s}$, and the last three channels of normalized quaternion $\mathbf{r}$.
Figure 17: More text-to-3D generation results (cont.). Our model is able to generate 2D Gaussian atlases to resemble high-quality 3D Gaussians from various prompts. We present the generated Gaussian atlases in the order from top left to bottom right: 3D location $\mathbf{x}$, albedo $\mathbf{c}$, color-coded opacity $\mathbf{o}$, normalized scale $\mathbf{s}$, and the last three channels of normalized quaternion $\mathbf{r}$.
D.2 More Generations
More results generated from a diverse list of prompts are shown in Figure 16 and Figure 17. We demonstrate impressive generation quality on text prompts involving descriptions for color, shape, style, abstract semantics, and quantities for objects. Additionally, we include videos of 360-degree renderings of the generated objects along with the supplementary material.
2D flattening and surface cutting may introduce unexpected artifacts in the final 3D generations [5]. However, when rendering the generated 2D atlases back as 3D Gaussians, no defects are observed at the ’seams’ of disconnected 2D boundaries, and our final 3D generations are coherent and natural, similar to other direct 3D generation approaches [10, 55]. This proves the concept of our approach — 3D generation achieved by a 2D diffusion model.
D.3 Diversity of Generations
With the same text prompt, we initiate the reverse diffusion with different random noise to demonstrate the diversity of generations in Figure 14. We observe a high diversity of geometries and appearances of the generated objects. This further proves that our proposed 2D representations of 3D Gaussians does not negatively alter the underlying semantics of the original 3D properties.
Appendix E Discussions
Limitations.
Our experiments revealed a trade-off between generation quality and the parameter $N$, which denotes the number of Gaussians per Gaussian atlas. A larger $N$ allows for finer-grained details in the generated outputs, but it also increases both training and inference costs. In this paper, we set $N=128\times 128=16,384$ Gaussians as a good compromise between computational efficiency and quality. However, in our current setup — using approximately three times fewer Gaussians than in [55] — our 3D generations do not exhibit significantly finer details compared to state-of-the-art methods that employ substantially more Gaussians.
Moreover, since this work repurposes a specific checkpoint of the LD model, no major updates have been made to the model architecture. Consequently, our model inherits the limitations of the original LD model for 3D generation, including reduced quality for long text prompts (see Figure 15).
Future Directions.
The purpose of this work is to introduce a novel way to represent 3D contents as 2D and make attempts to unify both 2D and 3D generation, allowing advancements in either paradigm to benefit both. There are two directions particularly worth pursuing by following the pathway presented in this work. The first is to experiment with more advanced diffusion models, such as transformer-based architectures [31] or latent diffusion models for Gaussians [35]. The second is to integrate plug-ins that are originaly designed for 2D diffusion models, such as IP-adapter [54] and DreamBooth [37], into 3D generations, given the unified diffusion framework.

### v2

@@
 Moreover, GVGen [10] and GaussianCube [57] achieve 3D Gaussian diffusion by transforming sparsely located Gaussians into more structured 3D volumes.
-GVGen fits 3D Gaussians directly on a volume with offsets, while GaussianCube applies Optimal Transport to move unorganized 3D Gaussians to vertices of a predefined 3D grid.
-Most recently, Gaussian Anything [17] adopts a two-step approach that generates Gaussians by auto-encoding latents in the point cloud space.
-In this work, we propose a novel structuralization approach that enables 2D transformation for 3D Gaussians, to leverage the power of well pretrained 2D diffusion models and unify the architecture of both 2D and 3D generation.
+GVGen fits 3D Gaussians directly on a volume with offsets, while GaussianCube applies Optimal Transport to move 3D Gaussians to vertices of a predefined 3D grid.
+Gaussian Anything [17] adopts a two-step approach that generates Gaussians by auto-encoding point cloud latents.
+Recently, TRELLIS [51] designed a structured representation that supports decoding to various 3D formats.
+In this work, we propose to transform 3D Gaussians onto 2D atlases, to leverage the power of well pretrained 2D diffusion models and unify the architecture of 2D and 3D generation.
 2D representations of 3D content.
@@
 Several works share motivations similar to ours.
-For example, PI3D [20] fine-tunes a pretrained text-to-image generation model to produce Triplanes represented as “pseudo-images,” while HexaGen3D [27] leverages pretrained image diffusion models for improved Triplane generation.
+For example, PI3D [21] fine-tunes a pretrained text-to-image generation model to produce Triplanes represented as “pseudo-images”, while HexaGen3D [28] leverages pretrained image diffusion models for improved Triplane generation.
 However, their fine-tuned diffusion models exhibit limited capacity, and an additional minutes-long refinement process is typically required to enhance the 3D generation.
 In addition to Triplane-based representations, Splatter Image [45] trains a 2D network to infer pixel-wise Gaussian attributes from single images; Omages [55] trains 2D diffusion models to fit 2D UV maps of the geometry and materials of 3D objects, ultimately reconstructing a textured 3D mesh.
-However, Omegas supports only class-conditioned generations of a limited pre-defined categories; GIMDiffusion [5] adopts a similar approach by representing the surface of 3D objects with geometry images [8] and leveraging 2D diffusion models in the UV space.
-However, their generated UV maps are sparse, leading to artifacts at the seams of the assembled 3D mesh.
-In contrast, our method introduces a dense 2D representation of 3D Gaussians, thereby mitigating these limitations.
+However, Omegas supports only class-conditioned generations of a limited pre-defined categories; GIMDiffusion [5] adopts a similar approach by representing the surface of 3D objects with geometry images [8] and leveraging 2D diffusion models in the UV space; DiffSplat [19] trains VAE to compress multi-view 3D Gaussians into 2D latents and achieve multi-view generation through a pretrained 2D diffusion model.
+In contrast, our method introduces a more dense 2D representation of 3D Gaussians, yielding a novel approach for repurposing pretrained 2D diffusion models.
 3 GaussianVerse
@@
 # Gaussians per fitting Total # fittings Compute time
-10,435$\pm$4,453 205,737 $>$3.8 GPU Years
+[…]
 Table 1: Statistics for the proposed GaussianVerse dataset.
@@
 See Figure 11 in the supplementary materials for our fitting results on objects with varying levels of detail and Figure 2 (d) for the distribution of the number of Gaussians across all fittings in GaussianVerse.
-With the above fitting strategy, we optimize per-object 3D Gaussians by minimizing photometric losses against multi-view RGB renderings:
+We optimize per-object 3D Gaussians by minimizing photometric losses against multi-view RGB renderings:
 $\lambda_{rgb}^{\prime}\mathcal{L}_{rgb}+\lambda_{ssim}^{\prime}\mathcal{L}_{ssim}+\lambda_{lpips}^{\prime}\mathcal{L}_{lpips}+\lambda_{reg}^{\prime}\mathcal{R},$ (2)
@@
 These disparities in data availability and representational complexity between 2D and 3D motivate our approach to leveraging learned priors from pre-trained 2D diffusion models for 3D Gaussian generation.
-However, sparse and unstructured Gaussians in 3D space cannot be directly passed to 2D models, which require inputs $\mathbf{X}$ to have: (i) only 2 spatial dimensions $W$ and $H$ such that $\mathbf{X}\in\mathbb{R}^{W\times H\times\cdot}$; (ii) valid “pixels” at each vertex of a dense 2D grid; and (iii) values within a specific distribution, either $\mathbf{X}\in[-1,1]$ for the VAE or $\mathbf{X}\sim\mathcal{N}(0,1)$ for the latent UNet.
+However, unstructured Gaussians in 3D space cannot be directly passed to 2D models, which require inputs $\mathbf{X}$ to have: (i) only 2 spatial dimensions; (ii) valid “pixels” at each vertex of a dense 2D grid; and (iii) values within a specific distribution, either $\mathbf{X}\in[-1,1]$ for the VAE or $\mathbf{X}\sim\mathcal{N}(0,1)$ for the denoiser.
 To make 3D Gaussians compatible with 2D diffusion models, we propose Gaussian Atlas, a 2D representation of 3D Gaussians.
@@
 Therefore, instead of exact mappings between 3D geometries and their 2D UVs, we focus on one characteristic of UV unwrapping — geometry flattening.
-Specifically, our goal is to find a mapping function $\mathcal{M}(\cdot)$ that transforms the 3D positions $\{\mathbf{x}\in\mathbb{R}^{3}\}$ of 3D Gaussians to 2D planar coordinates $\{\mathbf{\hat{x}}\in\mathbb{R}^{2}\}$: ($\mathcal{M}(\{\mathbf{x}\})\rightarrow\{\mathbf{\hat{x}}\}$.
+Specifically, our goal is to find a mapping function $\mathcal{M}(\cdot)$ that transforms the 3D positions $\{\mathbf{x}\in\mathbb{R}^{3}\}$ of 3D Gaussians to 2D planar coordinates $\{\mathbf{\hat{x}}\in\mathbb{R}^{2}\}$: $\mathcal{M}(\{\mathbf{x}\})\rightarrow\{\mathbf{\hat{x}}\}$.
 A natural approach is to parameterize such a function $\mathcal{M}$ as neural networks which can be optimized toward 2D and 3D consistency [58].
@@
 Our transportation process is also much faster due to the adaptive number of Gaussians.
-After positioning the 3D Gaussians on the surface of a unit sphere, we employ the classic equirectangular projection [49] as our function $\mathcal{M}$ to obtain flattened 2D coordinates $\{p_{i}\in\mathbb{R}^{2}\}$ of the 3D Gaussians.
-Lastly, in the plane offsetting stage, we apply another OT to map $\{p_{i}\}$ to the vertices $\{q_{i}\in\mathbb{R}^{2}\}$ of a square 2D grid with spatial size $\sqrt{N}\times\sqrt{N}$, further reducing sparsity.
-In particular, since now $\mathcal{M}$ is a deterministic function, the mapping between $\{p_{i}\}$ and $\{q_{i}\}$ remains identical between different 3D objects.
-This consistency allows us to perform OT in the last stage only once and reuse the computed indices for all objects.
-We refer to the final grid-like 2D representation of 3D Gaussians as Gaussian Atlas, with each atlas $\mathbf{X}$ in the shape of $\sqrt{N}\times\sqrt{N}\times(||\mathbf{x}-\mathbf{s}||+||\mathbf{c}||+||\mathbf{o}||+||\mathbf{s}||+||\mathbf{r}||)$ entailing all attributes of 3D Gaussians.
+[…]
 5 2D Diffusion for 3D Gaussian Generation
@@
 Our 3D generations exhibit the highest quality, minimal artifacts, and the best alignment with text prompts.
-In contrast, DreamGaussian [45], LGM [46], and TriplaneGaussian [58] struggle to produce natural and complete 3D assets, while GaussianCube [55] fails to capture key phrases in text prompts, such as “on a plate” and “pink rooftop.”
+In contrast, DreamGaussian [46], LGM [47], and TriplaneGaussian [60] struggle to produce natural and complete 3D assets, while GaussianCube [57] fails to capture key phrases in text prompts, such as “on a plate” and “pink rooftop”.
 3D Gaussians are characterized by five attributes: the 3D location (mean) $\mathbf{x}$, albedo $\mathbf{c}$, and scale $\mathbf{s}$, each represented as three-channel features; opacity $\mathbf{o}$, which is a single-channel feature; and rotation $\mathbf{r}$, typically represented as a four-channel quaternion.
@@
 $\lambda_{diff}\mathcal{L}_{diff}+\lambda_{rgb}\mathcal{L}_{rgb}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{lpips}\mathcal{L}_{lpips},$ (4)
-where $\mathcal{L}_{rgb},\mathcal{L}_{mask}$ are L1 loss on RGB renderings and accumulated opacity maps and $\mathcal{L}_{lpips}$ is the perceptual loss.
-$\lambda$s are the corresponding loss weights.
-The overall pipeline is outlined in Figure 4.
+[…]
 6 Experiments
-In this section, we benchmark and evaluate the effectiveness of our proposed methods on zero-shot text-to-3D generation, one of the most fundamental tasks in generative 3D modeling.
+In this section, we benchmark and evaluate our proposed methods on zero-shot text-to-3D generation, one of the most fundamental tasks in generative 3D modeling.
 Figure 6: Additional qualitative results.
@@
 This choice of $N$ accommodates the large variance of 3D Gaussians in GaussianVerse.
-We initialize the 2D UNet $\mathcal{F}$ from a LD checkpoint pretrained on LAION [39] and use CLIP [33] to encode text prompts.
+We initialize the 2D UNet $\mathcal{F}$ from a LD pretrain checkpoint [40] and use CLIP [34] to encode text prompts.
 The UNet is then fine-tuned with classifier-free guidance by randomly zeroing out the text encodings with a probability of 20%.
@@
 The LD UNet was finetuned for 1M steps on 8 A100 GPUs.
-For inference, we start from 2D random noise in the same shape as a Gaussian atlas and generate clean samples via reverse diffusion using the adaptive DPMsolver+++++ + [22] with a guidance scale of 3.5.
+For inference, we start from 2D random noise in the same shape as a Gaussian atlas and generate clean samples via reverse diffusion using the adaptive DPMsolver++++ [23] with a guidance scale of 3.5.
 A 3DGS sample can be generated and rendered in less than 5 seconds.
@@
 Qualitative results.
-We present zero-shot 3D generation results in Figure 5.
+We present text-to-3D generation results in Figure 5.
 Compared to DreamGaussian and LGM, our generations exhibit high visual fidelity without excessive details.
-In contrast to TriplaneGaussian and GaussianCube, our method aligns closely with the text prompts, reflecting more precise prompt conditioning in the generations.
-According to the additional qualitative results shown in Figure 6, our fine-tuned 2D diffusion model is capable of generating high-quality Gaussian atlases that effectively generalize to 3D information.
+In contrast to TriplaneGaussian and GaussianCube, our method aligns closely with the text prompts, reflecting more precise prompt conditioning.
+According to the additional results shown in Figure 6, our fine-tuned 2D diffusion model is capable of generating high-quality Gaussian atlases.
 Consequently, our generated 3D contents are coherent and self-contained, with no seams or artifacts that would typically be observed from UV unwrapping methods.
-We provide more qualitative results in the supplementary materials.
+Quantitative results.
+In Table 2, we present the CLIP and VQA scores for comparison methods.
+Our method outperforms all counterparts while generating the minimum number of Gaussians required.
+Specifically, it uses only $\frac{1}{2}$ the number of Gaussians and the number of training steps on 3D data compared to GaussianCube, achieving a CLIP score higher by 0.9 and a VQA score higher by 17%.
+This further supports our claim that pretrained 2D diffusion models can be repurposed for 3D content generation.
 User studies [^1] The user study was conducted entirely by researchers at Stanford..
-The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts.
+The user study was conducted with diverse participants who were asked to assess the overall generation quality to given text prompts.
 They performed pairwise comparisons between generations by our method and those from a competing method.
-We calculated the win rate from the collected $>2,500$ valid responses.
-As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube.
-The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
-User studies.
-The user study was conducted with a diverse group of participants who were asked to assess the overall generation quality corresponding to given text prompts.
-They performed pairwise comparisons between generations by our method and those from a competing method.
-We calculated the win rate from the collected $>2,500$ valid responses.
-As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to the state-of-the-art GaussianCube.
+We calculated the win rate from >2,500>2,500 valid responses.
+As shown in Figure 7, over 65% of users preferred the 3D content generated by our method when compared to GaussianCube.
 The preference rate was even more pronounced against TriplaneGaussian, with 88% of users favoring our method.
@@
 In this section, we conduct extensive studies and ablative experiments to validate our proposed approach from three key perspectives.
-$\bullet$  Learned prior knowledge in text-to-image diffusion models is universally transferable.
-Without pre-training on large-scale 2D datasets, the model struggles to develop a comprehensive understanding of natural language and content generation when trained solely on 3D data.
-As shown in Figure 8, we effectively repurpose the pre-trained 2D diffusion model, resulting in significantly faster convergence and higher-quality 3D generation.
+$\bullet$  Learned knowledge in text-to-image diffusion models is universally transferable.
+Without pre-training on large-scale 2D data, the model struggles to develop a comprehensive understanding of natural language and content generation when trained solely on 3D data.
+As shown in Figure 8, we effectively repurpose the pre-trained 2D diffusion model with faster convergence and higher-quality 3D generation.
 The qualitative results show that, compared to training from scratch, our approach produces more coherent and structured 3D assets earlier in training.
@@
 With capturable visual patterns, our proposed flattening strategy leads to faster model convergence and much higher generation quality at the same training step.
-Figure 9: The optimization-based flattening approach results in inconsistent patterns and poses challenges for fine-tuning.
-We present two objects (a) and their corresponding Gaussian atlases (3D location $\mathbf{x}$ only) obtained using our proposed flattening approach (b) and the optimization-based approach [56] (c), which generates only noisy Gaussians (d) and atlases (e) after training for the same number of steps.
+Figure 9: The optimization-based flattening approach results in inconsistent patterns that are hard for fine-tuning.
+We show two objects (a) and their Gaussian atlases (3D location $\mathbf{x}$ only) obtained using our proposed flattening approach (b) and the optimization-based approach [58] (c), which generates only noisy Gaussians (d) and atlases (e) after training for the same number of steps.
 $\bullet$  A pre-trained 2D diffusion model serves as a better initialization for 3D generation.
@@
 We observe overall small changes in the weights from the pre-trained UNet.
-Even at the layer with the greatest change, the difference is still 8×8$\times$8 × smaller than that from a random initialization.
+Even at the layer with the greatest change, the difference is still $8\times$ smaller than that from a random initialization.
 This indicates that the pre-trained LD weights provide better starting points for 3D generation, facilitating easier convergence to a local minimum.
@@
 We also provide extensive discussions and ablation studies that validate our claims and offer valuable insights into the effectiveness of our approach.
+Acknowledgments.
+Tiange Xiang, Scott Delp, Ehsan Adeli, and Li Fei-Fei’s work at Stanford were partially funded by the NIH Grant R01AG089169 and P41EB027060, Panasonic Holdings Corporation, the Gordon and Betty Moore Foundation, the Jaswa Innovator Award, Stanford HAI, Stanford HAI graduate fellowship, and Stanford Wu Tsai Human Performance Alliance.
+We thank Fei Jiang and Moustafa Meshry for their insightful feedback on this work.
 Supplementary Material
@@
 GaussianCube [57] NeurIPS 2024 125,653 32,768
-ShapeSplat [26] 3DV 2025 $\sim$65,000 $>$20,000
+ShapeSplat [27] 3DV 2025 $\sim$65,000 >>20,000
 GaussianVerse (ours) - 205,737 10,435